OpenStudy (anonymous):

solve the equations, y(y+2) = 3

OpenStudy (anonymous):

distribute the y first: y^2+2y = 3 this is equivalent to y^2+2y-3 = 0 when you factor, you get (y-1)(y+3) = 0 y = 1 and y = -3

OpenStudy (anonymous):

wait, I dont get it yet. how did you get (y-1)(y+3) ?

OpenStudy (anonymous):

when you factor. do you know how to factor?

OpenStudy (anonymous):